how would you write an equation for each line example: through (-2, 1) and parellel to Y = -3x+1?
Answer by venugopalramana(3286)

how would you write an equation for each line example: through (-2, 1) and parellel to Y = -3x+1?
if the second line is parallel to the first then write the same equation as the first replacing the constant term with say k which is to be found out
Y = -3X+k
if this is passing through -2,1 then this equation should be satisfied by -2,1
which helps us to find k
we get ...1 =-3*(-2)+k = 6+k
6+k=1
k=1-6 = -5
hence equation of a parallel line through (-2, 1) is
Y = -3X - 5